OpenStudy (anonymous):

simplify: (1/x + 1) / (1/x - 1)

OpenStudy (vishweshshrimali5):

Is this the question: \(\cfrac{\cfrac{1}{x} + 1}{\cfrac{1}{x} - 1}\)

OpenStudy (anonymous):

yes

OpenStudy (vishweshshrimali5):

Good now if x is not 0, then I can take x as LCM in both numerator and denominator. What will you get ?

OpenStudy (anonymous):

um I'm not quite sure

OpenStudy (vishweshshrimali5):

Ok let me help u there: \(\cfrac{\cfrac{1+x}{x}}{\cfrac{1-x}{x}}\)

OpenStudy (anonymous):

would you cancel out the x's?

OpenStudy (vishweshshrimali5):

Yes great. Now tell me what will you get?

OpenStudy (anonymous):

[1+x \div 1-x\]

OpenStudy (vishweshshrimali5):

Very nice.

OpenStudy (vishweshshrimali5):

Now that's your answer.

OpenStudy (anonymous):

Thank you!

OpenStudy (vishweshshrimali5):

Brilliant work and keep it up. Try to understand the basic concepts and you will clear every examination with flying colours. Bye and enjoy your time at OS. Gud day !

OpenStudy (anonymous):

How do you find the inverse of an equation?

OpenStudy (vishweshshrimali5):

Well actually you have to understand that inverse has various meanings in various contexts. So, you will have to specify it. Try to give me an example and I will help you out. :)

OpenStudy (anonymous):

\[f(x) = \frac{ 2x-3 }{ 5 }\]

OpenStudy (vishweshshrimali5):

Ok: See actually the best way to solve such type of question is this: (1) put f(x) = y (2) Find value of x in terms of y (3) Replace y by x and x by y.

OpenStudy (vishweshshrimali5):

See: (1) f(x) = y = (2x-3)/5 (2) Find value of x in terms of y: 5y = 2x-3 => 5y+3 = 2x => x = (5y+3)/2 (3) replace x by y and y by x So you get y = (5x+3)/2 Now put this y = g(x) So, g(x) = (5x+3)/2 Now, this g(x) will be the inverse function of f(x).

OpenStudy (vishweshshrimali5):

Did you get it?

OpenStudy (anonymous):

Yes thank you so much again!

OpenStudy (vishweshshrimali5):

Hey there is no problem. It was my pleasure.

OpenStudy (anonymous):

if an absolute value decreases doesn't the graph narrow?

OpenStudy (vishweshshrimali5):

Lets find out with the help of a graph: Lets say : |x| was initially 2 then it became less than 2. Okay ?

OpenStudy (anonymous):

Okay!

OpenStudy (vishweshshrimali5):

Now lets see its graph: Initially:  Finally:  (sorry couldn't shade it properly) Now the equation |x| < 2 signifies all values of x such that -2<x<2. So, we cannot say about narrowing of graph.